*Use a system of equations to solve (2 variables and 2 equations)*:
In order to complete her science experiment, Merrie needs 20 liters of a solution which is 25% alcohol. The only two alcohol solutions available in the school laboratory are a solution which is 40% alcohol and another which is 15% alcohol. How many liters of each of these two solutions must she mix together to produce the required solution?

Answer by stanbon(75887)

Use a system of equations to solve (2 variables and 2 equations)*:
In order to complete her science experiment, Merrie needs 20 liters of a solution which is 25% alcohol. The only two alcohol solutions available in the school laboratory are a solution which is 40% alcohol and another which is 15% alcohol. How many liters of each of these two solutions must she mix together to produce the required solution?
-----
Quantity: t + f = 20 liters
--------
Alcohol: 0.40t + 0.15f = 0.25*20
------------------------------------
Modify:
40t + 40f = 40*20
40t + 15f = 25*20
-----------------------
Subtract and solve for "f":
25f = 300
f = 12 liters (amt. of 15% solution needed)
----
t = 20-12 = 8 liters (amt. of 40% solution needed)
----------------
Cheers,
Stan H.
-----------------